Ari has been paying attention to the number of calories she burns while exercising. One day she spent three hours hiking and four hours canoeing, and she calculated that she burned 2 338 calories. The next day, she hiked for seven hours and canoed for one hour, and she calculated that she burned 3 322 calories. How many calories per hour does Ari burn doing each exercise?
Answer by Cromlix(4381)

Hi there,
Make hiking = x
Canoeing = y
3x + 4y = 2338.....(1)
7x + y = 3322......(2)
Multiply (2) by 4
3x + 4y = 2338.....(1)
28x + 4y = 13288......(2)
Subtract (1) from (2)
25x = 10950
x = 10950/25
x = 438
Substitute x = 438 into (1)
3x + 4y = 2338.....(1)
3(438) + 4y = 2338
1314 + 4y = 2338
4y = 2338 - 1314
4y = 1024
y = 256
Hiking = 438 calories
Canoeing = 256 calories.
Hope this helps :-)